### v2

# The Eisenbud–Goto conjecture for projectively normal varieties with mild singularities

Abstract.
For a nondegenerate projective variety $X$, the Eisenbud–Goto conjecture asserts that $\operatorname{reg}X\leq\operatorname{deg}X-\operatorname{codim}X+1$. Despite the existence of counterexamples, identifying the classes of varieties for which the conjecture holds remains a major open problem. In this paper, we prove that the Eisenbud–Goto conjecture holds for $2$-very ample projectively normal varieties with factorial, rational, hypersurface singularities and isolated Gorenstein singularities.
Key words and phrases:
Eisenbud–Goto conjecture, Castelnuovo–Mumford regularity, syzygies, projectively normal varieties, mild singularities
2020 Mathematics Subject Classification:
Primary 14N05; Secondary 13D02
1. Introduction
In this paper, we work over an algebraically closed field $\mathbb{K}$ of characteristic zero and require varieties to be integral. Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate projective variety. For a coherent sheaf $\mathcal{F}$ on $\mathbb{P}^{r}$, we say $\mathcal{F}$ is $m$-regular if
$H^{i}(\mathbb{P}^{r},\mathcal{F}(m-i))=0$
for all $i\geq 1$. The regularity $\operatorname{reg}\mathcal{F}$ of $\mathcal{F}$ is defined as the least integer $m$ such that $\mathcal{F}$ is $m$-regular. The regularity $\operatorname{reg}X$ of $X$ is defined as the regularity of its ideal sheaf $\mathcal{I}_{X}$. Regularity is one of the invariants measuring the complexity of projective varieties. In 1984, D. Eisenbud and S. Goto conjectured a bound on the regularity of projective varieties.
Conjecture 1.1 (The Eisenbud–Goto conjecture, [10]).
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate projective variety. Then
$\operatorname{reg}X\leq\deg X-\operatorname{codim}X+1.$ (1)
It holds for arithmetically Cohen–Macaulay varieties. In 1983, L. Gruson, R. Lazarsfeld, and C. Peskine [14] showed the conjecture for curves. For the case of smooth surfaces, H. C. Pinkham [38] and R. Lazarsfeld [24] proved the conjecture. In 1990, Z. Ran [39] asserted that the conjecture holds for most smooth threefolds of codimension at least 6 using local differential geometric methods. In 1998, I. Peeva and B. Sturmfels [37] proved the conjecture for projective toric varieties of codimension two. In 1998, S. Kwak [22] proved the weak bound
$\operatorname{reg}X\leq\deg X-\operatorname{codim}X+2$ (2)
for smooth threefolds and $\operatorname{reg}X\leq\deg X-\operatorname{codim}X+5$ for smooth fourfolds. In 1999, S. Kwak [23] showed (1) for smooth threefolds in $\mathbb{P}^{5}$. For smooth projective varieties of dimension at most 14, L. Chiantini, N. Chiarli, and S. Greco [7] showed
$\operatorname{reg}X\leq\deg X-\operatorname{codim}X+1+\frac{(n-2)(n-1)}{2},$
hence improving the bound for smooth fourfolds to $\operatorname{reg}X\leq\deg X-\operatorname{codim}X+4$. Here, the condition $\dim X\leq 14$ comes from Mather’s theorem [27]. In 2015, W. Niu [33] showed (1) for normal surfaces with rational, Gorenstein elliptic, log canonical singularities. In 2022, W. Niu and J. Park [32] showed the weak bound (2) for normal threefolds with rational singularities.
In 2018, J. McCullough and I. Peeva [29] published a surprising result that there exist counterexamples to the Eisenbud–Goto conjecture. Their construction uses Rees-like algebras and step-by-step homogenization. In particular, they constructed a singular threefold of degree 31 and regularity 38 in $\mathbb{P}^{5}$. Furthermore, they showed there does not exist a polynomial function of $\deg X$ bounding $\operatorname{reg}X$. In 2024, the author and S. Kwak [15] constructed surface counterexamples using binomial rational maps between projective spaces. In 2025, J. Choe [9] constructed counterexamples using the unprojection method.
Although counterexamples exist, it is still a very interesting problem to identify the classes of varieties for which the conjecture holds. As J. McCullough and I. Peeva stated in [30], the following are interesting cases where the conjecture may hold:
  1. (1)
    smooth projective varieties;
  2. (2)
    projectively normal varieties;
  3. (3)
    projective toric varieties; and
  4. (4)
    projective singular surfaces.
Case (4) was ruled out in [15].
On the other hand, the validity of the Eisenbud–Goto conjecture for $X$ is equivalent to the following two conditions:
  1. (1)
    for all $m\geq\deg X-\operatorname{codim}X$, the variety $X\subseteq\mathbb{P}^{r}$ is $m$-normal, i.e., $H^{0}(\mathbb{P}^{r},\mathcal{O}_{\mathbb{P}^{r}}(m))\to H^{0}(X,\mathcal{O}_{X}(m))$ is surjective; and
  2. (2)
    $\operatorname{reg}\mathcal{O}_{X}\leq\deg X-\operatorname{codim}X$.
In 2014, A. Noma [35] showed
$\operatorname{reg}\mathcal{O}_{X}\leq\deg X-\operatorname{codim}X$ (3)
for smooth projective varieties except when $X$ is a scroll over a smooth projective curve. Here, a scroll over a smooth curve $C$ is defined as the image of $\mathbb{P}_{C}(\mathcal{E})$ for some locally free sheaf $\mathcal{E}$ over $C$ under a birational morphism defined by a sublinear system of $\lvert\mathcal{O}_{\mathbb{P}_{C}(\mathcal{E})}(1)\rvert$. In 2020, S. Kwak and J. Park [21] showed (3) for smooth scrolls over smooth projective curves. Therefore (3) holds for smooth projective varieties. This implies that the Eisenbud–Goto conjecture holds for smooth projectively normal varieties. In 2021, J. Moraga, J. Park, and L. Song [31] showed (3) for normal projective varieties with isolated $\mathbb{Q}$-Gorenstein singularities. Hence the Eisenbud–Goto conjecture holds for projectively normal varieties with isolated $\mathbb{Q}$-Gorenstein singularities.
In this paper, we prove that the Eisenbud–Goto conjecture holds for 2-very ample projectively normal varieties with mild singularities. Regarding mild singularities, we allow two types of singularities: one is the factorial, rational, hypersurface singularity and the other is the isolated Gorenstein singularity. Specifically, we prove the following.
Theorem A.
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate $2$-very ample projectively normal variety with factorial, rational, hypersurface singularities and isolated Gorenstein singularities. Then
$\operatorname{reg}X\leq\deg X-\operatorname{codim}X+1.$
This paper is organized as follows. In Section 2, we review some known results. In Section 3, we discuss vanishing theorems for singular varieties. In Section 4, we derive the birational double point formula for factorial points. In Section 5, we prove the main theorem.
Acknowledgement
The author was supported by a KIAS Individual Grant (MG101401) at Korea Institute for Advanced Study. The author thanks Professor Sijong Kwak for his encouragement.
2. Preliminaries
In this section, we collect some known results for later discussion. Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate projective variety of dimension $n$ and codimension $e$. We recall the definition of $k$-very ampleness, as the main theorem relies on this positivity condition.
Definition 2.1.
Let $\mathcal{L}$ be a line bundle on $X$. We say $\mathcal{L}$ is $k$-very ample if
$H^{0}(X,\mathcal{L})\to H^{0}(\xi,\mathcal{L}|_{\xi})$
is surjective for any subscheme $\xi$ of dimension 0 and length $k+1$. When $X\subseteq\mathbb{P}^{r}$ is embedded by the complete linear system $\lvert\mathcal{L}\rvert$, we say $X$ is $k$-very ample if $\mathcal{L}$ is $k$-very ample.
On locally factorial varieties, reflexive sheaves of rank 1 are invertible. This fact is used in the proof of Lemma 4.3.
Proposition 2.2 ([18, Proposition 1.9]).
On integral Noetherian locally factorial schemes, any reflexive sheaf of rank 1 is invertible.
The Zariski–Fujita theorem states that if the base locus of a line bundle is finite, then the line bundle is semiample.
Theorem 2.3 ([40], also cf. [25, Remark 2.1.31]).
Let $\mathcal{L}$ be a line bundle on $X$ with
$\dim\operatorname{Bs}\lvert\mathcal{L}\rvert\leq 0.$
Then $\mathcal{L}$ is semiample.
Now we review the results by A. Noma. First, we denote the locus of smooth points of $X$ that induce a nonbirational inner projection to $\mathbb{P}^{r-1}$ by
$\mathcal{C}(X):=\{x\in X_{\textnormal{sm}}\mid\operatorname{length}(X\cap\langle x,y\rangle)\geq 3\text{ for general }y\in X\}$
where $\langle\cdot\rangle$ denotes the linear span.
It is worth noting that the closure of this locus is a union of linear subspaces.
Theorem 2.4 ([34, Theorem 5]).
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate projective variety of codimension $e\geq 2$ and $\Lambda$ be an irreducible component of $\mathcal{C}(X)$. Then $\overline{\Lambda}$ is linear and
$\dim\mathcal{C}(X)\leq\min\{n-1,\dim X_{\operatorname{sing}}+2\}.$
Let $x_{1},\cdots,x_{m}$ be general points in $X$ and $Z:=\{x_{1},\cdots,x_{m}\}$ where $1\leq m\leq e-1$. Denote by $\pi:X\setminus\langle Z\rangle\to\mathbb{P}^{r-m}$ the linear projection of $X\subseteq\mathbb{P}^{r}$ from $\langle Z\rangle$ and the closure of the image of $\pi$ by $X_{Z}$. The following theorem shows the map $\pi^{\prime}:X\setminus\langle Z\rangle\to X_{Z}$ called the induced projection is isomorphic at $x$ if $x\in X_{\textnormal{sm}}\setminus\mathcal{C}(X)$. Note that we pick general points $x_{1},\cdots,x_{m}$ when $x$ is given, so one can not expect that $\pi^{\prime}$ is isomorphic on the entire $X_{\textnormal{sm}}\setminus\mathcal{C}(X)$.
Theorem 2.5 ([35, Lemma 2.2]).
Assume as above and $e\geq 2$. Suppose $x\in X_{\textnormal{sm}}\setminus\mathcal{C}(X)$. Then $\pi^{\prime}:X\setminus\langle Z\rangle\to X_{Z}$ is isomorphic at $x$.
Denote by $E_{x_{1},\cdots,x_{m}}(X)$ the union of positive dimensional fibers of $\pi$, i.e.,
$E_{x_{1},\cdots,x_{m}}(X):=\{x\in X\setminus\langle Z\rangle\mid\dim\pi^{-1}(\pi(x))\geq 1\}.$
Definition 2.6.
We say $X$ satisfies $(E_{m})$ if
$\dim E_{x_{1},\cdots,x_{m}}(X)\geq n-1,$
or equivalently, $\dim E_{x_{1},\cdots,x_{m}}(X)=n-1$, where $x_{1},\cdots,x_{m}\in X$ are general points ($1\leq m\leq e-1$).
Let $\operatorname{Bl}_{Z}X$ be the blowing up of $X$ along $Z$. Then we get the map $\tilde{\pi}:\operatorname{Bl}_{Z}X\to\mathbb{P}^{r-m}$ called the extended projection. With this notation, the following lemma holds.
Lemma 2.7 ([35, Lemma 1.3]).
Assume as above and $n\geq 2$. Then
$\tilde{E}_{x_{1},\cdots,x_{m}}(X):=\{x\in\operatorname{Bl}_{Z}X\mid\dim\tilde{\pi}^{-1}(\tilde{\pi}(x))\geq 1\}$
is closed in $\operatorname{Bl}_{Z}X$ and $\dim E_{x_{1},\cdots,x_{m}}(X)\geq n-1$ if and only if $\dim\tilde{E}_{x_{1},\cdots,x_{m}}(X)\geq n-1$.
Using the condition $(E_{m})$, Noma established the positivity of the double point divisor for inner projections.
Theorem 2.8 ([35, Theorem 1]).
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate smooth projective variety of dimension $n\geq 2$, codimension $e\geq 2$, and degree $d$ not satisfying $(E_{m})$ where $1\leq m\leq e-1$. Then
$\operatorname{Bs}\lvert\mathcal{O}_{X}(d-n-m-2)\otimes\omega_{X}^{-1}\rvert\subseteq\mathcal{C}(X)$
and $\mathcal{O}_{X}(d-n-m-1)\otimes\omega_{X}^{-1}$ is very ample on $X\setminus\mathcal{C}(X)$. Furthermore if $\dim\mathcal{C}(X)\leq 0$, then $\mathcal{O}_{X}(d-n-m-2)\otimes\omega_{X}^{-1}$ is semiample so that $\mathcal{O}_{X}(d-n-m-1)\otimes\omega_{X}^{-1}$ is ample.
We define a scroll over a smooth curve as follows.
Definition 2.9.
A scroll over a smooth curve $C$ is the image of $\mathbb{P}_{C}(\mathcal{E})$ for some locally free sheaf $\mathcal{E}$ over $C$ under a birational morphism defined by a sublinear system of $\lvert\mathcal{O}_{\mathbb{P}_{C}(\mathcal{E})}(1)\rvert$.
Noma classified projective varieties satisfying $(E_{m})$ for $1\leq m\leq e-1$.
Theorem 2.10 ([35, Theorem 3]).
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate projective variety of dimension $n\geq 2$ and codimension $e$.
  1. (1)
    When $e\geq 2$ and $X$ satisfies $(E_{1})$, $X$ is projectively equivalent to a scroll over a smooth curve. If furthermore $X$ is smooth, $X$ is projectively equivalent to $\mathbb{P}_{C}(\mathcal{E})\subseteq\mathbb{P}^{r}$ where $\mathcal{E}$ is a locally free sheaf on a smooth projective curve.
  2. (2)
    When $e\geq 3$ and $X$ satisfies $(E_{2})$ but not $(E_{1})$, $X$ is projectively equivalent to a cone over the Veronese surface $\nu_{2}(\mathbb{P}^{2})\subseteq\mathbb{P}^{5}$. If furthermore $X$ is smooth, $X$ is projectively equivalent to the Veronese surface $\nu_{2}(\mathbb{P}^{2})\subseteq\mathbb{P}^{5}$.
  3. (3)
    When $e\geq 4$, $X$ satisfies $(E_{m})$ for some $3\leq m\leq e-1$ if and only if $X$ satisfies $(E_{1})$.
In particular, if $n\geq 2$ and $X$ is not projectively equivalent to
  1. (1)
    a scroll over a smooth curve $C$; and
  2. (2)
    a cone over the Veronese surface $\nu_{2}(\mathbb{P}^{2})\subseteq\mathbb{P}^{5}$,
then $X$ does not satisfy $(E_{m})$ for any $1\leq m\leq e-1$.
3. Vanishing theorems for singular varieties
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate projective variety of dimension $n$. We denote the normalized dualizing complex of $X$ by $\omega_{X}^{\bullet}$ (cf. [16]) and define the dualizing sheaf as $\omega_{X}:=\mathcal{H}^{-n}(\omega_{X}^{\bullet})$, the lowest cohomology of $\omega_{X}^{\bullet}$. When $X$ is Cohen–Macaulay, this is the only cohomology that does not vanish. D. Arapura, D. B. Jaffe, and L. Song [1, 2] studied Kodaira vanishing theorem for singular projective varieties. One of their vanishing results for singular projective varieties is the following.
Theorem 3.1 ([2, Lemma 3.3]).
Let $X\subseteq\mathbb{P}^{r}$ be a projective variety and $\mathcal{L}$ be a nef and big line bundle on $X$. Then
$H^{i}(X,\mathcal{H}^{j}(\omega_{X}^{\bullet})\otimes\mathcal{L})=0$
for any positive integer $i>\dim X_{\operatorname{sing}}$ and any $j\in\mathbb{Z}$.
The key part of the proof is to utilize the following vanishing theorem which is a special case of [12, Théorème 3.1.A] or a consequence of Grauert–Riemenschneider vanishing and Kawamata–Viehweg vanishing.
Theorem 3.2 (cf. [12, Théorème 3.1.A]).
Let $X\subseteq\mathbb{P}^{r}$ be a projective variety, $\mathcal{L}$ a nef and big line bundle on $X$, and $f:Y\to X$ be a resolution of singularities. Then
$H^{i}(X,f_{*}\omega_{Y}\otimes\mathcal{L})=0$
for all $i\geq 1$.
We define the locus
$X_{\operatorname{CM}}:=\{x\in X\mid\text{$X$ is Cohen--Macaulay at $x$}\}.$
This locus is open in $X$ (cf. [28, Exercise 24.2]). We also define the locus
[…]
This locus is open in $X$ [11].
With these loci, one can express the vanishing theorem for singular varieties (Theorem 3.1) as follows. Lemma 3.3 is used in the proof of Theorem 5.6.
Lemma 3.3.
Let $X\subseteq\mathbb{P}^{r}$ be a projective variety of dimension $n$ and $\mathcal{L}$ be a nef and big line bundle on $X$. Then
$H^{i}(X,\mathcal{H}^{j}(\omega_{X}^{\bullet})\otimes\mathcal{L})=0$
for any nonnegative integer $i>\dim(X\setminus X_{\operatorname{CM}})$ when $j\geq-n+1$ and for any positive integer $i>\dim(X\setminus X_{\operatorname{rat}})$ when j=nj=-n. In particular, it holds for any positive integer $i>\dim(X\setminus X_{\operatorname{rat}})$ and any $j\in\mathbb{Z}$.
Proof.
Denote $n:=\dim X$. For $j\geq-n+1$, the sheaf $\mathcal{H}^{j}(\omega_{X}^{\bullet})$ vanishes on $X_{\operatorname{CM}}$. Hence it is supported on $X\setminus X_{\operatorname{CM}}$ so that
$H^{i}(X,\mathcal{H}^{j}(\omega_{X}^{\bullet})\otimes\mathcal{L})=0$
for any nonnegative integer $i>\dim(X\setminus X_{\operatorname{CM}})$ and any integer $j\geq-n+1$.
Now let j=nj=-n. For a resolution of singularities $f:Y\to X$, we have the injective trace map $\operatorname{Trace}_{Y/X}:f_{*}\omega_{Y}\to\omega_{X}$ (cf. [20, Proposition 5.77]). Also,
$H^{i}(X,f_{*}\omega_{Y}\otimes\mathcal{L})=0$
for all $i>0$ (Theorem 3.2). On the other hand, the cokernel of $\operatorname{Trace}_{Y/X}$ is supported on $X\setminus X_{\operatorname{rat}}$. Hence $H^{i}(X,\omega_{X}\otimes\mathcal{L})=0$ for all positive $i>\dim(X\setminus X_{\operatorname{rat}})$. The last statement follows from $X\setminus X_{\operatorname{CM}}\subseteq X\setminus X_{\operatorname{rat}}$. ∎
4. The birational double point formula
The purpose of this section is to prove Lemma 4.3, which is used in the proof of Theorem 5.6. R. Lazarsfeld [26] and A. Noma [35] stated the birational double point formula which is derived from the double point formula [13, Theorem 9.3].
Lemma 4.1 ([26, Lemma 10.2.8] and [35, Lemma 2.1]).
Let $f:X\to P$ be a morphism between smooth projective varieties where $\dim X=n$ and $\dim P=n+1$. Suppose $f$ factors through a surjective birational morphism $g:X\twoheadrightarrow Y$ for a subvariety $Y\subseteq P$ of codimension 1. Then for some effective divisors $D$ and $E$ on $X$, it holds that
$D-E\equiv-K_{X}+f^{*}(K_{P}+Y)$
and $E$ is exceptional for $g$. Furthermore, if $f$ is isomorphic at a point $x\in X$, then $x\notin\operatorname{Supp}(D-E)$.
In [31], the authors derived the birational double point formula as follows.
Lemma 4.2 ([31, Lemma 2.3]).
Let $f:X\to\mathbb{P}^{n+1}$ be a morphism from a normal projective variety of dimension $n$. Suppose $f$ factors through a surjective birational morphism $g:X\twoheadrightarrow Y$ for a hypersurface $Y\subseteq\mathbb{P}^{n+1}$. Then for some effective Weil divisors $D$ and $E$ on $X$, it holds that
$D-E\equiv-K_{X}+f^{*}(K_{P}+Y)$
and $E$ is exceptional for $g$. Furthermore, if $f$ is isomorphic at a smooth point $x\in X_{\textnormal{sm}}$, then $x\notin\operatorname{Supp}(D-E)$.
Here, $E$ is exceptional for $g$ means that the image of every irreducible component of $E$ has dimension lower than $E$. The divisor $D$ is called the double point divisor. The positivity of the double point divisor has been investigated in several works [3, 19, 35, 8].
In the following, we derive the birational double point formula for factorial points to prove Theorem 5.6. To state it, we define the locus
$X_{\operatorname{fac}}:=\{x\in X\mid\text{$X$ is factorial at $x$}\}.$
This locus is open in $X$ [4], also cf. [5].
Lemma 4.3.
Let $f:X\to Y$ be a birational morphism between Gorenstein projective varieties of dimension $n$ not admitting any exceptional divisor. Suppose $X$ is normal. If $f$ is isomorphic at a factorial point $x\in X_{\operatorname{fac}}$, then $\lvert\omega_{X}^{-1}\otimes f^{*}\omega_{Y}\rvert\neq\varnothing$ and $x\notin\operatorname{Bs}\lvert\omega_{X}^{-1}\otimes f^{*}\omega_{Y}\rvert$.
Proof.
Let $i:X_{\operatorname{fac}}\to X$ be the immersion. Denote $g:=f\circ i$ so that we have the following commutative diagram.
[…]
Note that $g$ is birational and does not admit an exceptional divisor. By Proposition 2.2, the sheaf $(g^{*}g_{*}\omega_{X_{\operatorname{fac}}})^{\vee\vee}$ is a line bundle since it is reflexive of rank 1 and $X_{\operatorname{fac}}$ is locally factorial. The counit map $g^{*}g_{*}\omega_{X_{\operatorname{fac}}}\to\omega_{X_{\operatorname{fac}}}$ factors through $(g^{*}g_{*}\omega_{X_{\operatorname{fac}}})^{\vee\vee}$. The map $(g^{*}g_{*}\omega_{X_{\operatorname{fac}}})^{\vee\vee}\to\omega_{X_{\operatorname{fac}}}$ is injective since $g$ is birational and $(g^{*}g_{*}\omega_{X_{\operatorname{fac}}})^{\vee\vee}$ is torsion free. Hence $(g^{*}g_{*}\omega_{X_{\operatorname{fac}}})^{\vee\vee}=\omega_{X_{\operatorname{fac}}}(-E)$ for some effective divisor $E$ that is exceptional for $g$. As $g$ does not admit an exceptional divisor, we get
$(g^{*}g_{*}\omega_{X_{\operatorname{fac}}})^{\vee\vee}=\omega_{X_{\operatorname{fac}}}.$
On the other hand, as $f$ is a birational morphism between projective varieties, there exists the trace map $f_{*}\omega_{X}\to\omega_{Y}$ that is injective. Note that $i_{*}\omega_{X_{\operatorname{fac}}}=i_{*}i^{*}\omega_{X}=\omega_{X}$ as $X$ is normal and $\dim(X\setminus X_{\operatorname{fac}})\leq n-2$. Hence
$g_{*}\omega_{X_{\operatorname{fac}}}\otimes\omega_{Y}^{-1}=f_{*}i_{*}\omega_{X_{\operatorname{fac}}}\otimes\omega_{Y}^{-1}=f_{*}\omega_{X}\otimes\omega_{Y}^{-1}\hookrightarrow\mathcal{O}_{Y}.$
This induces the map
$\omega_{X_{\operatorname{fac}}}\otimes g^{*}\omega_{Y}^{-1}=(g^{*}g_{*}\omega_{X_{\operatorname{fac}}})^{\vee\vee}\otimes g^{*}\omega_{Y}^{-1}=(g^{*}g_{*}\omega_{X_{\operatorname{fac}}}\otimes g^{*}\omega_{Y}^{-1})^{\vee\vee}\to(g^{*}\mathcal{O}_{Y})^{\vee\vee}=\mathcal{O}_{X_{\operatorname{fac}}}$
which is injective as $g$ is birational. Therefore $\omega_{X_{\operatorname{fac}}}\otimes g^{*}\omega_{Y}^{-1}$ is an invertible ideal sheaf on $X_{\operatorname{fac}}$ so that
$\omega_{X_{\operatorname{fac}}}\otimes g^{*}\omega_{Y}^{-1}=\mathcal{O}_{X_{\operatorname{fac}}}(-D_{0})$
for some effective divisor $D_{0}$.
As $X$ is normal, we have $\dim(X\setminus X_{\operatorname{fac}})\leq n-2$ so that the divisor $D_{0}$ extends to an effective Weil divisor $D$ on $X$. Also note that $i_{*}(\omega_{X_{\operatorname{fac}}}^{-1})=i_{*}i^{*}(\omega_{X}^{-1})=\omega_{X}^{-1}$. Applying $i_{*}$ to $\mathcal{O}_{X_{\operatorname{fac}}}(D_{0})=\omega_{X_{\operatorname{fac}}}^{-1}\otimes g^{*}\omega_{Y}$, we get
$\mathcal{O}_{X}(D)=i_{*}(\omega_{X_{\operatorname{fac}}}^{-1}\otimes i^{*}f^{*}\omega_{Y})=i_{*}(\omega_{X_{\operatorname{fac}}}^{-1})\otimes f^{*}\omega_{Y}=\omega_{X}^{-1}\otimes f^{*}\omega_{Y}.$
Hence $D\in\lvert\omega_{X}^{-1}\otimes f^{*}\omega_{Y}\rvert$.
Suppose $f$ is isomorphic at a factorial point $x\in X_{\operatorname{fac}}$. Then the above injective map
$\omega_{X_{\operatorname{fac}}}\otimes g^{*}\omega_{Y}^{-1}\to\mathcal{O}_{X_{\operatorname{fac}}}$
is isomorphic at $x$ so that $x\notin D_{0}$. As $x\in X_{\operatorname{fac}}$, this implies $x\notin D$. Therefore $x\notin\operatorname{Bs}\lvert\omega_{X}^{-1}\otimes f^{*}\omega_{Y}\rvert$. ∎
5. The Eisenbud–Goto regularity bound
In [35], Noma bounded the regularity of $\mathcal{O}_{X}$ using his results on the positivity of the double point divisor (Theorem 2.8), the classification of projective varieties satisfying $(E_{m})$ (Theorem 2.10), and the characterization of smooth projective varieties with $\dim\mathcal{C}(X)\geq 1$ ([36, Corollary 6.2]).
Theorem 5.1 ([35, Corollary 5]).
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate smooth projective variety of dimension $n\geq 2$, codimension $e$, and degree $d$ that is not projectively equivalent to a scroll over a smooth projective curve. Then
$\operatorname{reg}\mathcal{O}_{X}\leq d-e.$
The remaining case was solved by S. Kwak and J. Park [21].
Theorem 5.2 ([21, Proposition 3.6]).
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate smooth scroll over a smooth projective curve $C$ of dimension $n\geq 2$, codimension $e$, and degree $d$. Then
  1. (1)
    $\operatorname{reg}\mathcal{O}_{X}\leq d-e$;
  2. (2)
    $\operatorname{reg}\mathcal{O}_{X}=d-e$ if and only if $X$ is a rational normal scroll; and
  3. (3)
    $\operatorname{reg}\mathcal{O}_{X}=d-e-1$ if and only if $X$ is an isomorphic projection of rational normal scroll from a point or an elliptic normal surface scroll with $d=e+3$.
Furthermore, J. Moraga, J. Park, and L. Song [31] derived the result for normal projective varieties with isolated $\mathbb{Q}$-Gorenstein singularities.
Theorem 5.3 ([31, Theorem 1]).
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate normal variety with isolated $\mathbb{Q}$-Gorenstein singularities of codimension $e$ and degree $d$. Then
$\operatorname{reg}\mathcal{O}_{X}\leq d-e.$
To prove the main theorem, we introduce two lemmas: Lemma 5.4 and Lemma 5.5. These lemmas follow the argument of [35, Lemma 2.2] (Theorem 2.5) and extend the statement to singular points $x$ with $\dim T_{x}X\leq n+1$.
Lemma 5.4.
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate projective variety of dimension $n$ and codimension $e\geq 2$. For a point $x\in X$, suppose that $\dim T_{x}X\leq n+1$ and $\operatorname{length}(X\cap\langle x,x_{1}\rangle)=2$ for the general point $x_{1}\in X$. Let $X_{x_{1}}$ be the closure of the image of the linear projection $\pi:X\setminus\{x_{1}\}\to\mathbb{P}^{r-1}$ of $X$ from $x_{1}$. Then the induced projection $\pi^{\prime}:X\setminus\{x_{1}\}\to X_{x_{1}}$ is isomorphic at $x$. Furthermore, if $e\geq 3$, then $\operatorname{length}(X_{x_{1}}\cap\langle\pi(x),y_{\operatorname{gen}}\rangle)=2$ for the general point $y_{\operatorname{gen}}\in X_{x_{1}}$.
Proof.
Note that $x$ is not a vertex of $X$ by the assumption. Let $b:\operatorname{Bl}_{x_{1}}X\to X$ be the blowing up of $X$ along $x_{1}$ with the exceptional divisor $E_{1}$ and $\tilde{x}$ be the preimage of $x$ via $b$. By the generic smoothness theorem (cf. [17, Lemma III.10.5]), the inner projection of $X$ from $x$ is unramified at $x_{1}$, so we have $x\notin T_{x_{1}}X$. Thus the extended projection $\operatorname{Bl}_{x_{1}}X\to\mathbb{P}^{r-1}$ separates $\tilde{x}$ from any point in $E_{1}$. Also as $\operatorname{length}(X\cap\langle x,x_{1}\rangle)=2$, the extended projection $\operatorname{Bl}_{x_{1}}X\to\mathbb{P}^{r-1}$ separates $\tilde{x}$ from any other point in $\operatorname{Bl}_{x_{1}}X\setminus E_{1}$. Therefore $\operatorname{Bl}_{x_{1}}X\to\mathbb{P}^{r-1}$ separates $\tilde{x}$ from any other point in $\operatorname{Bl}_{x_{1}}X$. Furthermore, $x_{1}\notin T_{x}X$ by the assumption that $\dim T_{x}X\leq n+1$ and $e\geq 2$. Therefore $\operatorname{Bl}_{x_{1}}X\to\mathbb{P}^{r-1}$ is an embedding near $\tilde{x}$, so $\pi^{\prime}:X\setminus\{x_{1}\}\to X_{x_{1}}$ is isomorphic at $x$.
Now we assume $e\geq 3$ to show the last statement. It is enough to show $\operatorname{length}(X_{x_{1}}\cap\langle\pi(x),\pi(y)\rangle)=2$ for the general point $y\in X$. We suppose the contrary, i.e.,
$\operatorname{length}(X_{x_{1}}\cap\langle\pi(x),\pi(y)\rangle)\geq 3$
for the general point $y\in X$ or $\pi(x)$ is a vertex point of $X_{x_{1}}$. The latter case cannot occur since $\dim T_{\pi(x)}X_{x_{1}}\leq n+1$ as $\pi^{\prime}:X\setminus\{x_{1}\}\to X_{x_{1}}$ is isomorphic at $x$.
We claim that there exists a point $z\in X\setminus\{x_{1}\}$ such that $\pi(z)\in\langle\pi(x),\pi(y)\rangle$, $\pi(z)\neq\pi(x)$, and $\pi(z)\neq\pi(y)$. As $\dim T_{\pi(x)}X_{x_{1}}\leq n+1$, we have $\dim\langle T_{\pi(x)}X_{x_{1}},x_{1}\rangle\leq n+2$. As $e\geq 3$, this implies $y\notin\langle T_{\pi(x)}X_{x_{1}},x_{1}\rangle$ so that $\pi(y)\notin T_{\pi(x)}X_{x_{1}}$. Also, by the generic smoothness theorem, the inner projection of $X_{x_{1}}$ from $\pi(x)$ is unramified at $\pi(y)$, so we have $\pi(x)\notin T_{\pi(y)}X_{x_{1}}$. Furthermore, as $\dim\langle T_{x_{1}}X,x\rangle=n+1$, we have $y\notin\langle T_{x_{1}}X,x\rangle$ so that any points in $E_{1}$ do not map to $\langle\pi(x),\pi(y)\rangle$. Therefore the claim holds.
By the claim, the points $x_{1},x,y,z$ are distinct points on a $2$-plane. Furthermore, $z\notin\langle x,x_{1}\rangle$ and $z\notin\langle x,y\rangle$ as
$\operatorname{length}(X\cap\langle x,x_{1}\rangle)=\operatorname{length}(X\cap\langle x,y\rangle)=2$
by the assumption. Let $X_{x}$ be the closure of the image of the inner projection $\pi_{x}:X\setminus\{x\}\to\mathbb{P}^{r-1}$ of $X$ from $x$. Then $\langle\pi_{x}(x_{1}),\pi_{x}(y)\rangle$ is the general secant line to $X_{x}$, but is also a trisecant line as it contains $\pi_{x}(z)$. This contradicts the general position lemma (cf. [35, Lemma 1.2]). ∎
Lemma 5.5.
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate projective variety of dimension $n$ and codimension $e\geq 2$. For a point $x\in X$, suppose that $\dim T_{x}X\leq n+1$ and $\operatorname{length}(X\cap\langle x,x_{\operatorname{gen}}\rangle)=2$ for the general point $x_{\operatorname{gen}}\in X$. Let $x_{1},\cdots,x_{e-1}\in X$ be general points and $Z:=\{x_{1},\cdots,x_{e-1}\}$. Denote by $X_{Z}$ the closure of the image of the linear projection $\pi:X\setminus\langle Z\rangle\to\mathbb{P}^{n+1}$ of $X$ from $\langle Z\rangle$. Then the induced projection $\pi^{\prime}:X\setminus\langle Z\rangle\to X_{Z}$ is isomorphic at $x$.
Proof.
This follows by applying Lemma 5.4 inductively. Note that when $e\geq 3$, Lemma 5.4 ensures that the condition
$\operatorname{length}(X\cap\langle x,x_{\operatorname{gen}}\rangle)=2\text{ for the general point }x_{\operatorname{gen}}\in X$ (4)
holds also in the next step. When $e=2$, Lemma 5.4 does not ensure that (4) holds for the image, but it holds for $X$ itself, which suffices to apply the projection. ∎
In the following, the proof follows the argument of Noma (Theorem 2.8) which considers smooth projective varieties. We extend it to the singular case under consideration.
Theorem 5.6.
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate $2$-very ample normal Gorenstein variety. Suppose that
  1. (1)
    $X$ has factorial rational singularities; and
  2. (2)
    $\dim T_{p}X\leq\dim X+1$ for $p\in X$
possibly except some finite points. Then
$\operatorname{reg}\mathcal{O}_{X}\leq\deg X-\operatorname{codim}X.$
Proof.
Denote $n:=\dim X$, $e:=\operatorname{codim}X$, and $d:=\deg X$. We assume $n\geq 2$ and $e\geq 2$ since the case $n=1$ is well known [14] and the case $e=1$ is trivial. By the assumption, $X$ is neither a cone nor a scroll over a smooth curve. If $X$ is the Veronese surface $\nu_{2}(\mathbb{P}^{2})$, then the result is well known as it has minimal degree. Thus we may assume $X$ is not $\nu_{2}(\mathbb{P}^{2})$. Then $X$ does not satisfy ($E_{e-1}$) by Theorem 2.10.
Let $Y$ be a cofinite subset of $X_{\operatorname{sing}}$ such that
  1. (1)
    $X$ has factorial rational singularities on $Y$; and
  2. (2)
    $\dim T_{p}X\leq n+1$ for all $p\in Y$.
Such $Y$ must exist by the assumption. Pick a point $x\in X_{\textnormal{sm}}\cup Y$ and e1e-1 general points $x_{1},\cdots,x_{e-1}$ in $X$. Denote $Z:=\{x_{1},\cdots,x_{e-1}\}$. Let $b:\operatorname{Bl}_{Z}X\to X$ be the blowing up of $X$ along $Z$ with exceptional divisors $E_{1},\cdots,E_{e-1}$ such that each $E_{i}$ corresponds to $x_{i}$. Note that $b$ is isomorphic at $x$. We denote the preimage of $x$ by $\tilde{x}$, i.e., $b(\tilde{x})=x$. Let $\pi:X\setminus\langle Z\rangle\to\mathbb{P}^{n+1}$ be the linear projection of $X$ from the linear span $\langle Z\rangle$ of $Z$. We denote the closure of the image of $\pi:X\setminus\langle Z\rangle\to\mathbb{P}^{n+1}$ by $X_{Z}\subseteq\mathbb{P}^{n+1}$ and denote the induced projection by $\pi^{\prime}:X\setminus\langle Z\rangle\to X_{Z}$. Also, let $\tilde{\pi}:\operatorname{Bl}_{Z}X\to\mathbb{P}^{n+1}$ be the extended projection and $\tilde{\pi}^{\prime}:\operatorname{Bl}_{Z}X\to X_{Z}$ be the induced morphism.
[…]
The morphism $\tilde{\pi}^{\prime}$ is birational as $\pi^{\prime}$ is. The birational map $\pi^{\prime}$ does not admit an exceptional divisor as $X$ does not satisfy $(E_{e-1})$. Then $\tilde{\pi}^{\prime}$ also does not admit an exceptional divisor by Lemma 2.7.
The projection $\pi^{\prime}:X\setminus\langle Z\rangle\to X_{Z}$ is isomorphic at $x$ by Lemma 5.5. Also, by the generic smoothness theorem, we have $x\notin T_{x_{1}}X=\langle T_{x_{1}}X,x_{1}\rangle$. As $\dim\langle x,T_{x_{1}}X,x_{1}\rangle=n+1$, we have $x_{2}\notin\langle x,T_{x_{1}}X,x_{1}\rangle$. Thus $x\notin\langle T_{x_{1}}X,x_{1},x_{2}\rangle$. Repeating this inductively, we get $x\notin\langle T_{x_{1}}X,x_{1},\cdots,x_{e-1}\rangle$. This shows that $x\notin\langle T_{x_{i}}X,x_{1},\cdots,x_{e-1}\rangle$ for all $i=1,\cdots,e-1$. Hence $\tilde{\pi}^{\prime}:\operatorname{Bl}_{Z}X\to X_{Z}$ separates $\tilde{x}$ from any point in $E:=E_{1}+\cdots+E_{e-1}$. Together with the fact that $\pi^{\prime}:X\setminus\langle Z\rangle\to X_{Z}$ is isomorphic at $x$, this implies that $\tilde{\pi}^{\prime}:\operatorname{Bl}_{Z}X\to X_{Z}$ is isomorphic at $\tilde{x}$.
On the other hand, as $X_{Z}$ is a hypersurface of $\mathbb{P}^{n+1}$, it is Gorenstein. Also, $\operatorname{Bl}_{Z}X$ is a normal Gorenstein variety as $X$ is a normal Gorenstein variety and $x_{1},\cdots,x_{e-1}$ are smooth points. As $x\in X_{\textnormal{sm}}\cup Y$, it holds that $X$ is factorial at $x$ so that $\operatorname{Bl}_{Z}X$ is factorial at $\tilde{x}$. Hence by Lemma 4.3, $\tilde{x}\notin\operatorname{Bs}\lvert\omega_{\operatorname{Bl}_{Z}X}^{-1}\otimes(\tilde{\pi}^{\prime})^{*}\omega_{X_{Z}}\rvert$. Let $\tilde{D}\in\lvert\omega_{\operatorname{Bl}_{Z}X}^{-1}\otimes(\tilde{\pi}^{\prime})^{*}\omega_{X_{Z}}\rvert$ be the divisor not containing $\tilde{x}$. As $b:\operatorname{Bl}_{Z}X\to X$ is isomorphic outside $E:=E_{1}+\cdots+E_{e-1}$, by taking the isomorphic image of $\tilde{D}|_{\operatorname{Bl}_{Z}X\setminus E}$ and taking the closure, we get an effective divisor $D$ on $X$ not containing $x$ such that
$\mathcal{O}_{X}(D)|_{X\setminus Z}=(\omega_{X}^{-1}\otimes(\pi^{\prime})^{*}\omega_{X_{Z}})|_{X\setminus Z}.$
This equality extends to the entire $X$ since $X$ is normal with $n\geq 2$, the set $Z$ is finite, and $(\omega_{X}^{-1}\otimes(\pi^{\prime})^{*}\omega_{X_{Z}})|_{X\setminus Z}$ is a line bundle on $X\setminus Z$. Note that $\deg X_{Z}=d-e+1$. Hence
$\omega_{X_{Z}}=(\omega_{\mathbb{P}^{n+1}}\otimes\mathcal{O}_{\mathbb{P}^{n+1}}(X_{Z}))|_{X_{Z}}=\mathcal{O}_{X_{Z}}(d-n-e-1),$
so $D\in\lvert\omega_{X}^{-1}(d-n-e-1)\rvert$. Therefore $x\notin\operatorname{Bs}\lvert\omega_{X}^{-1}(d-n-e-1)\rvert$. By repeating this for all $x\in X_{\textnormal{sm}}\cup Y$, we get
$\operatorname{Bs}\lvert\omega_{X}^{-1}(d-n-e-1)\rvert\subseteq X_{\operatorname{sing}}\setminus Y.$
Hence $\omega_{X}^{-1}(d-n-e-1)$ is semiample by Zariski–Fujita’s theorem (Theorem 2.3). Therefore $\omega_{X}^{-1}(d-n-e)$ is ample. Note that $\dim(X\setminus X_{\operatorname{rat}})\leq 0$ by the assumption. Hence by Lemma 3.3, we get
$H^{i}(X,\mathcal{O}_{X}(j))=0\text{ for all $1\leq i\leq n$ and $j\geq d-n-e$.}$
Thus $\mathcal{O}_{X}$ is (de)(d-e)-regular. ∎
We obtain A as a corollary.
Theorem A.
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate $2$-very ample projectively normal variety with factorial, rational, hypersurface singularities and isolated Gorenstein singularities. Then
$\operatorname{reg}X\leq\deg X-\operatorname{codim}X+1.$
Proof.
As before, denote $e:=\operatorname{codim}X$ and $d:=\deg X$. Since any factorial Cohen–Macaulay point is a Gorenstein point (cf. [6, Corollary 3.3.19]), $X$ is Gorenstein. Hence by Theorem 5.6, $\mathcal{O}_{X}$ is (de)(d-e)-regular. From the long exact sequence
$\cdots\to H^{i}(\mathbb{P}^{r},\mathcal{O}_{\mathbb{P}^{r}}(j))\to H^{i}(X,\mathcal{O}_{X}(j))\to H^{i+1}(\mathbb{P}^{r},\mathcal{I}_{X}(j))\to H^{i+1}(\mathbb{P}^{r},\mathcal{O}_{\mathbb{P}^{r}}(j))\to\cdots,$
it follows that
$H^{i+1}(\mathbb{P}^{r},\mathcal{I}_{X}(d-e-i))=0\text{ for all }i\geq 1.$
As $X$ is projectively normal, we have $H^{1}(\mathbb{P}^{r},\mathcal{I}_{X}(j))=0$ for all $j\geq 0$. Therefore, $\mathcal{I}_{X}$ is (de+1)(d-e+1)-regular. ∎